OpenStudy (anonymous):

What acceleration will a force of 20 newtons cause if applied to a go-kart with a mass of 20 kilograms

OpenStudy (anonymous):

Force = Mass times Acceleration \[F = MA\] Rearranging to solve for Acceleration. Acceleration = Force divided by Mass \[A=\frac{ F }{ M }\] Units must be the same Grams, meters, seconds. You can see you have Kilograms so a conversion is needed.

OpenStudy (anonymous):

Just noticed Kg is actually the proper units.

OpenStudy (anonymous):

if you are in S.I Units for Force. \[Newton = \frac{ Kg \times m }{ s^2 }\] \[Mass = Kg\] \[Acceleration = \frac{ m }{ s^2 }\] So you do not need a conversion just the formula to solve for A, acceleration.